# How individual vs shared coordination governs the degree of correlation in rotational vs residence times in a high-viscosity lithium electrolyte

### v1

@@ Abstract
Commercially used carbonate-based electrolytes in lithium-ion batteries are susceptible to many challenges, including flammability, volatility, and lower thermal stability. Solvated ionic liquids of LiTFSI salt (lithium bis(trifluoromethylsulfonyl)-amide) and glyme-based solvents are potential alternative candidates for commonly used electrolytes. We perform classical molecular dynamics (MD) simulations study the effect of concentration and temperature on the translational and rotational dynamics. The radial distribution function shows stronger coordination of Li+ ions with tetraglyme(G4), as shown in earlier studies, and forms a stable [Li(G4)]+ cation complex. The self-diffusion coefficients are lower than the values experimentally observed but show better improvement over other classical force fields. An increase in the salt concentrations leads to a higher viscosity of the system and reduces the overall ionic mobility of Li+ ions. Diluting the system with a larger number of glyme molecules leads to shorter rotational relaxation times for both TFSI and tetraglyme. Ion-residence times show that Li+ ions form stable and long-lasting complexes with G4 molecules than TFSI anions. The residence time of [Li(G4)]+ complex decreases at higher salt concentrations due to the availability of fewer G4 molecules to coordinate with a Li+ ion. G4 is also seen to form polydentate complexes with Li+ without a shared coordination, allowing rotation without breaking coordination, unlike TFSI, which requires coordination disruption for rotation. This distinction explains the poor correlation between rotation and residence time for G4 and the strong correlation for TFSI.
@@ I Introduction
This study focuses on investigating an extreme range of LiTFSI and G4 mixture compositions (from 1:4 to 2:1) at various temperatures with classical MD simulations using a modified OPLS force field.Jorgensen, Maxwell, and Tirado-Rives (1996) The partial charges for atoms of LiTFSI and G4 are reparametrized here (Fig. 1). A correct structure prediction and long timescale simulations performed in the present work provide a reliable set of results for diffusion coefficients, residence times, rotational relaxation times, cluster sizes and theorize a not yet explored interplay of these.
@@ II Computational Details
Four mixtures with different LiTFSI/G4 ratios (1:4, 1:2, 1:1, and 1:4) were studied, as shown in Table S1 of supplementary information. These mixtures were modeled using a charge-reparametrized OPLS force field. Partial atomic charges for the LiTFSI and G4 molecules were calculated (Fig. 1a) using a range-separated hybrid functional wB97x-D4 functionalChai and Head-Gordon (2008); Caldeweyher, Bannwarth, and Grimme (2017) with sufficiently large basis set def2-tzvpp, utilizing the ORCA 5.0.1 packageNeese (2022). The PMEEssmann et al. (1995) algorithm was used to compute the long-range electrostatic interactions. The cut-off for Lennard-Jones and electrostatic interactions is set to 10 Å. A time-step of 1 femtosecond is used throughout the simulations. All simulations were conducted using the Gromacs 2021 Abraham et al. (2015) software package. For post-processing and analysis, VMD 1.9.3 Humphrey, Dalke, and Schulten (1996), TravisBrehm et al. (2020), and in-house scripts developed in python were used.
@@ III.6 Ionic aggregates and clusters
It is important to notice here that G4 can form individual unshared complexes with a Li+ ion in a polydentate manner, which TFSI cannot. Thus, a rotation of Li…G4 complex can occur without the need of breaking the coordination, while in case of Li…TFSI, break in coordination is essential for any rotation of an individual TFSI. This explains the poor correlation between and for G4, and highly correlated those for TFSI. A longer residence time only restrict the rotation when there is coordination from multiple similar entities. A polydentate anion or solvent coordinating in an unshared manner would not affect its rotation with its residence time and thus can live longer while rotating faster.

### v2

@@ Abstract
Commercially used carbonate-based electrolytes in lithium-ion batteries are susceptible to many challenges, including flammability, volatility, and lower thermal stability. Solvated ionic liquids of LiTFSI salt (lithium bis(trifluoromethylsulfonyl)-amide) and glyme-based solvents are potential alternative candidates for commonly used electrolytes. We perform classical molecular dynamics (MD) simulations study the effect of concentration and temperature on the translational and rotational dynamics. The radial distribution function shows stronger coordination of Li+ ions with tetraglyme(G4), as shown in earlier studies, and forms a stable [Li(G4)]+ cation complex. The self-diffusion coefficients are lower than the values experimentally observed but show better improvement over other classical force fields. An increase in the salt concentrations leads to a higher viscosity of the system and reduces the overall ionic mobility of Li+ ions. Diluting the system with a larger number of glyme molecules leads to shorter rotational relaxation times for both TFSI and tetraglyme. Ion-residence times show that Li+ ions form stable and long-lasting complexes with G4 molecules than TFSI anions. The residence time of [Li(G4)]+ complex increases at higher salt concentrations due to the availability of fewer G4 molecules to coordinate with a Li+ ion. G4 is also seen to form polydentate complexes with Li+ without a shared coordination, allowing rotation without breaking coordination, unlike TFSI, which requires coordination disruption for rotation. This distinction explains the poor correlation between rotation and residence time for G4 and the strong correlation for TFSI.
@@ I Introduction
This study investigates the extreme range of LiTFSI and G4 mixture compositions (from 1:4 to 2:1) at various temperatures with classical MD simulations using a modified OPLS force-field.Jorgensen, Maxwell, and Tirado-Rives (1996) The partial charges for atoms of LiTFSI and G4 are reparametrized here (Fig. 1). A correct structure prediction and long timescale simulations performed in the present work provide a reliable set of results for diffusion coefficients, residence times, rotational relaxation times, and cluster sizes and theorize a not-yet-explored interplay of these.
@@ II Computational Details
Four mixtures with different LiTFSI/G4 ratios (1:4, 1:2, 1:1, and 1:4) were studied, as shown in Table S1 of supplementary information. These mixtures were modeled using a charge-reparameterized OPLS force field. Partial atomic charges for the LiTFSI and G4 molecules were calculated (Fig. 1a) using a range-separated hybrid functional wB97x-D4 functionalChai and Head-Gordon (2008); Caldeweyher, Bannwarth, and Grimme (2017) with a sufficiently large basis set def2-tzvpp, utilizing the ORCA 5.0.1 packageNeese (2022). The PMEEssmann et al. (1995) algorithm was used to compute the long-range electrostatic interactions. The cut-off for Lennard-Jones and electrostatic interactions is set to 10 Å. A time-step of 1 femtosecond is used throughout the simulations. All simulations were conducted using the Gromacs 2021 Abraham et al. (2015) software package. For post-processing and analysis, VMD 1.9.3 Humphrey, Dalke, and Schulten (1996), TravisBrehm et al. (2020), and in-house scripts developed in python were used.
@@ III.6 Ionic aggregates and clusters
It is important to notice here that G4 can form individual unshared complexes with a Li+ ion in a polydentate manner, which TFSI cannot. Thus, a rotation of Li…G4 complex can occur without the need of breaking the coordination, while in case of Li…TFSI, break in coordination is essential for any rotation of an individual TFSI. This explains the poor correlation between and for G4, and highly correlated those for TFSI. A longer residence time only restrict the rotation when there is coordination from multiple similar entities. A polydentate anion or solvent coordinating in an unshared manner would not affect its rotation with its residence time and thus can live longer while rotating faster. The correlation observed here is potentially an important descriptor, that can be further tested in extending the general theories for concentrated electrolytesKrucker-Velasquez and Swan (2021); Dinpajooh et al. (2024) to those with long-chain entities.